# Fully distributed consensus control for stochastic multi-agent systems under undirected and directed topologies

### 14 Mar 2026 v1

@@ I Introduction
A key development in this field is the widespread adoption of adaptive gain design, which enables dynamic control adjustment using only local information. Recent studies demonstrate this approach across diverse network topologies and system models. For instance, in [12], the consensus problem of second-order MASs with heterogeneous unknown time-varying inertias and control gains was studied under directed graphs through adaptive $\sigma$-modification schemes. [13] solved the problem of fully distributed consensus control for MASs with general linear dynamics under undirected graphs via adaptive dynamic event-triggered communication strategies. Further extending to signed networks, [14] proposes a time-varying parametric Lyapunov-based adaptive protocol for prescribed-time bipartite synchronization in cooperative-antagonistic networks. Studies such as [15] and [16] respectively ensure practical fixed-time consensus for single-integrator systems under undirected graphs and resolve leaderless/leader-following consensus for disturbed second-order systems under directed graphs, both employing distributed adaptive mechanisms. Moreover, [17] advances further by transcending topological and model constraints, proposing a topology-agnostic, data-driven event-triggered adaptive learning algorithm that achieves cooperative control under model uncertainty and communication constraints.
Collectively, these works exemplify the core idea of utilizing localized adaptive mechanisms to effectively operate across diverse network structures, ranging from undirected and directed graphs to signed digraphs, and progressing toward topology-agnostic scenarios, thereby illustrating an evolutionary trajectory from reliance on specific structural information to generalization independent of network topology. However, It should be noted that the aforementioned protocols primarily focus on deterministic MASs, leaving the challenge of stochastic disturbances largely unaddressed in current fully distributed control frameworks.
@@
iii) By employing advanced stochastic analysis tools including the functional $\mathrm{It\hat{o}}$ formula and stopping-time techniques, we rigorously establish the existence and uniqueness of solutions for the closed-loop stochastic MASs (path-dependent stochastic systems), which feature inherent high nonlinearity under our proposed fully distributed protocols.
@@ IV-A Fully distributed consensus control for stochastic MASs with the directed topology
Let $\xi(t)=[\xi_{1}^{T}(t),...,\xi_{N}^{T}(t)]^{T}$ and $\tilde{\sigma}(t)=\mathrm{diag}\{\sigma_{1}(t),...,\sigma_{N}(t)\}$. According to Lemma IV.1, the states $\xi(t)$ can be partitioned into $\xi_{l}(t)=[\xi_{1}^{T}(t),\dots,\xi_{M}^{T}(t)]^{T}$ and $\xi_{f}(t)=[\xi_{M+1}^{T}(t),\dots,\xi_{N}^{T}(t)]^{T}$, the states $\vartheta(t)$ can be partitioned into $\vartheta_{l}(t)=[\vartheta_{1}^{T}(t),\dots,\vartheta_{M}^{T}(t)]^{T}$ and $\vartheta_{f}(t)=[\vartheta_{M+1}^{T}(t),\dots,\vartheta_{N}^{T}(t)]^{T}$, the time-varying adaptive gain $\tilde{C}(t)$ can be partitioned into $\tilde{C}_{l}(t)=\mathrm{diag}\{c_{1}(t),...,c_{M}(t)\}$ and $\tilde{C}_{f}(t)=\mathrm{diag}\{c_{M+1}(t),...,c_{N}(t)\}$, the auxiliary time-varying gain $\tilde{\Sigma}(t)$ can be partitioned into $\tilde{\Sigma}_{l}(t)=\mathrm{diag}\{\Sigma_{1}(t),...,\Sigma_{M}(t)\}$ and $\tilde{\Sigma}_{f}(t)=\mathrm{diag}\{\Sigma_{M+1}(t),...,\Sigma_{N}(t)\}$, $\tilde{\sigma}(t)$ can be partitioned into $\tilde{\sigma}_{l}(t)=\mathrm{diag}\{\sigma_{1}(t),...,\sigma_{M}(t)\}$ and $\tilde{\sigma}_{f}(t)=\mathrm{diag}\{\sigma_{M+1}(t),...,\sigma_{N}(t)\}$, corresponding to the two subsets $\mathcal{V}_{l}$ and $\mathcal{V}_{f}$. Based on the closed-loop stochastic system (4), we obtain $d\vartheta_{l}(t)=(I_{M}\otimes A+(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t)\mathcal{L}_{11})\otimes B\mathcal{K})\vartheta_{l}(t)dt+(I_{M}\otimes C)\vartheta_{l}(t)dw(t)$, and $d\vartheta_{f}(t)=(I_{N-M}\otimes A+(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t)\mathcal{L}_{22})\otimes B\mathcal{K})\vartheta_{f}(t)dt+(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t)\mathcal{L}_{21}\otimes B\mathcal{K})\vartheta_{l}(t)dt+(I_{N-M}\otimes C)\vartheta_{f}(t)dw(t)$. Note that $\xi_{l}(t)=(\mathcal{L}_{11}\otimes I_{n})\vartheta_{l}(t)$ and $\xi_{f}(t)=(\mathcal{L}_{21}\otimes I_{n})\vartheta_{l}(t)+(\mathcal{L}_{22}\otimes I_{n})\vartheta_{f}(t)$. Then, we can get
@@
with $V_{l}(\vartheta_{t},t)=\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds+\eta_{2}^{l}\sum_{i=1}^{M}r_{i}(c_{i}(t)-\psi_{i})^{2}$ and $V_{f}(\vartheta_{t},t)=\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds+\eta_{2}^{f}\sum_{i=M+1}^{N}s_{i-M}(c_{i}(t)-\psi_{i})^{2}$, where $\eta_{1}^{l},\eta_{2}^{l},\eta_{1}^{f},\eta_{2}^{f}\geq 0$, $r_{i},s_{i}>0$, and $\psi_{i}$ are positive constants to be determined. According to the functional $\mathrm{It\hat{o}}$ formula and condition (I), it can be deduced that $dV(\vartheta_{t},t)$ has the following form
$\begin{split}dV&(\vartheta_{t},t)\!=\!\mathfrak{L}V(\vartheta_{t},t)dt+\frac{\partial V(\vartheta_{t},t)}{\partial\vartheta}(I_{N}\!\otimes\!C)\vartheta(t)dw(t),\end{split}$
where $\mathfrak{L}V(\vartheta_{t},t)$ is defined as
$\begin{split}\mathfrak{L}V(\vartheta_{t},t)=\mathfrak{L}V_{l}(\vartheta_{t},t)+\mathfrak{L}V_{f}(\vartheta_{t},t)\end{split}$ (12)
@@
with $\mathfrak{L}V_{l}(\vartheta_{t},t)=\eta_{1}^{l}\sum_{i=1}^{M}r_{i}\dot{c}_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds+2\eta_{2}^{l}\sum_{i=1}^{M}r_{i}(c_{i}(t)-\psi_{i})\dot{c}_{i}(t)+\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)-\int_{0}^{\sigma_{i}(t)}\mu k_{1}\frac{s}{c_{i}^{2}(t)}\dot{c}_{i}(t)(k_{2}+\frac{s}{{c}_{i}(t)})^{\mu-1}ds$, and $\mathfrak{L}V_{f}(\vartheta_{t},t)=\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}\dot{c}_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds+2\eta_{2}^{f}\sum_{i=M+1}^{N}s_{i-M}(c_{i}(t)-\psi_{i})\dot{c}_{i}(t)+\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)-\int_{0}^{\sigma_{i}(t)}\mu k_{1}\frac{s}{c_{i}^{2}(t)}\dot{c}_{i}(t)(k_{2}+\frac{s}{{c}_{i}(t)})^{\mu-1}ds$. Since $\mu,k_{1},k_{2}>0,\sigma_{i}(t)\geq 0$, we have $\mathfrak{L}V_{l}(\vartheta_{t},t)\leq\sum_{i=1}^{M}r_{i}\big[\eta_{1}^{l}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{l}c_{i}(t)-2\eta_{2}^{l}\psi_{i}\big]\dot{c}_{i}(t)+\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)$, and $\mathfrak{L}V_{f}(\vartheta_{t},t)\leq\sum_{i=M+1}^{N}s_{i-M}\big[\eta_{1}^{f}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{f}c_{i}(t)-2\eta_{2}^{f}\psi_{i}\big]\dot{c}_{i}(t)+\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)$. Let $\tilde{\psi}=\mathrm{diag}\{\psi_{1},...,\psi_{N}\}$. According to Lemma IV.1, $\tilde{\psi}$ can be partitioned into $\tilde{\psi}_{l}=\mathrm{diag}\{\psi_{1},...,\psi_{M}\}$ and $\tilde{\psi}_{f}=\mathrm{diag}\{\psi_{M+1},...,\psi_{N}\}$. Let $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$. Note that $\sum_{i=1}^{M}\sigma_{i}(t)=\xi_{l}^{T}(t)(I_{M}\otimes P)\xi_{l}(t)$ and $\xi_{l}(t)=(\mathcal{L}_{11}\otimes I_{n})\vartheta_{l}(t)$. By the definition of $\sigma_{i}(t)$, (10), (11) and using Lemma IV.1 and IV.2, we have
@@
$\begin{split}\mathfrak{L}&V(\vartheta_{t},t)\\ \leq&\eta_{1}^{l}\xi_{l}^{T}\!(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &\!\!\!\!\!+\!\eta_{1}^{f}\xi_{f}^{T}\!(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &+\xi_{l}^{T}\!(t)\!\Big[\Phi_{t}^{l}\otimes\!\Gamma\Big]\!\xi_{l}(t)+\xi_{f}^{T}\!(t)\!\Big[\Phi_{t}^{f}\otimes\!\Gamma\Big]\!\xi_{f}(t),\end{split}$ (17)
@@
$\begin{split}\mathfrak{L}&V(\vartheta_{t},t)\\ \leq&\eta_{1}^{l}\xi_{l}^{T}\!(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &+\!\eta_{1}^{f}\xi_{f}^{T}\!(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &-\xi_{l}^{T}\!(t)\Big[\big(\eta_{3}^{l}(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{2}\!+\!\check{\psi}_{l}\big)\otimes\Gamma\Big]\xi_{l}(t)\\ &-\xi_{f}^{T}\!(t)\Big[\big(\eta_{3}^{f}(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))^{2}\!+\!\check{\psi}_{f}\big)\otimes\Gamma\Big]\xi_{f}(t),\\ \leq&\eta_{1}^{l}\xi_{l}^{T}\!(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes(A^{T}P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &-\xi_{l}^{T}(t)\left[2\sqrt{\eta_{3}^{l}\check{\psi}_{l}}(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes\Gamma\right]\xi_{l}(t)\\ &+\eta_{1}^{f}\xi_{f}^{T}\!(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes(A^{T}P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &-\xi_{f}^{T}(t)\left[2\sqrt{\eta_{3}^{f}\check{\psi}_{f}}(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes\Gamma\right]\xi_{f}(t).\end{split}$
@@
$\begin{split}\mathfrak{L}V(\vartheta_{t},t)\leq&-\eta_{1}^{l}\xi_{l}^{T}(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes I_{n}]\xi_{l}(t)\\ &-\eta_{1}^{f}\xi_{f}^{T}(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes I_{n}]\xi_{f}(t)\\ \leq&-\eta_{1}^{l}\varepsilon\xi_{l}^{T}(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes P]\xi_{l}(t)\\ &-\eta_{1}^{f}\varepsilon\xi_{f}^{T}(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes P]\xi_{f}(t)\\ =&-\eta_{1}^{l}\varepsilon\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\sigma_{i}(t)\\ &-\eta_{1}^{f}\varepsilon\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\sigma_{i}(t)\\ \leq&-\eta_{1}^{l}\varepsilon\sum_{i=1}^{M}r_{i}c_{i}(t)\!\!\int_{0}^{\sigma_{i}(t)}\!\!\!k_{1}\Big(k_{2}+\frac{s}{c_{i}(t)}\Big)^{\mu}ds\\ &-\eta_{1}^{f}\varepsilon\!\!\!\sum_{i=M+1}^{N}\!\!\!\!\!s_{i-M}c_{i}(t)\!\!\!\int_{0}^{\sigma_{i}(t)}\!\!\!\!\!\!k_{1}\Big(k_{2}\!+\!\frac{s}{c_{i}(t)}\Big)^{\mu}ds\\ \leq&-\varepsilon V(\vartheta_{t},t)+\varphi(t),\end{split}$ (20)
@@
Step 2: By the functional $\mathrm{It\hat{o}}$ formula for $V(\vartheta_{t},t)$ and the estimation (20) of $\mathfrak{L}V(\vartheta_{t},t)\leq-\eta_{1}^{l}\varepsilon\xi_{l}^{T}(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes P]\xi_{l}(t)-\eta_{1}^{f}\varepsilon\xi_{f}^{T}(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes P]\xi_{f}(t)\leq-\eta_{1}\varepsilon\xi^{T}(t)[\tilde{S}(\tilde{C}(t)\tilde{\Sigma}(t))\otimes P]\xi(t)$ with $\eta_{1}=\eta_{1}^{l}\wedge\eta_{1}^{f}$ and $\tilde{S}=\mathrm{diag}\{R,S\}$, it can be deduced that
$\begin{split}&V(\vartheta_{t},t)-V(\vartheta_{0},0)\\ &=\int_{0}^{t}\mathfrak{L}V(\vartheta_{s},s)ds+M_{v1}(t)\\ &\leq-\eta_{1}\varepsilon\!\int_{0}^{t}\!\xi^{T}(s)[\tilde{S}(\tilde{C}(s)\tilde{\Sigma}(s))\!\otimes\!P]\xi(s)ds+M_{v1}(t),\end{split}$ (21)
@@
$\begin{split}\lim_{t\to\infty}\sup V(\vartheta_{t},t)<\infty\ a.s.\end{split}$
@@
$\begin{split}\ \mathrm{E}\!\!\int_{0}^{t}\!\!\eta_{1}\varepsilon\xi^{T}\!(s)[\tilde{S}(\tilde{C}(s)\tilde{\Sigma}(s))\!\otimes\!P]\xi(s)ds\leq\!\mathrm{E}V(\vartheta_{0},0)\!<\!\infty,\end{split}$

### 31 Mar 2026 v2

@@ I Introduction
A key development in this field is the widespread adoption of adaptive gain design, which enables dynamic control adjustment using only local information. Recent studies demonstrate this approach across diverse network topologies and system models. For instance, in [12], the consensus problem of second-order MASs with heterogeneous unknown time-varying inertias and control gains was studied under directed graphs through adaptive $\sigma$-modification schemes. In [13], the problem of fully distributed consensus for MASs with general linear dynamics under undirected graphs was solved via adaptive dynamic event-triggered communication strategies. Further extending to signed networks, a time-varying parametric Lyapunov-based adaptive protocol was proposed in [14] for prescribed-time bipartite synchronization in cooperative-antagonistic networks. Meanwhile, distributed adaptive mechanisms were employed in [15] and [16] to achieve practical fixed-time consensus for single-integrator systems under undirected graphs and leaderless/leader-following consensus for second-order systems under directed graphs, respectively. Moreover, a topology-agnostic, data-driven event-triggered adaptive learning algorithm was introduced in [17] for cooperative control under model uncertainty and communication constraints. Collectively, these works demonstrate that the adaptive mechanisms can effectively operate across diverse network structures, ranging from undirected and directed graphs to signed digraphs and topology-agnostic scenarios. However, it should be noted that the aforementioned protocols primarily focus on deterministic MASs, leaving the challenge of stochastic disturbances largely unaddressed in current fully distributed control frameworks.
@@
iii) By employing stochastic analysis tools including the functional $\mathrm{It\hat{o}}$ formula and stopping-time techniques, we rigorously establish the existence and uniqueness of solutions for a class of highly nonlinear, path-dependent stochastic systems.
@@ IV-A Fully distributed consensus control for stochastic MASs with the directed topology
Let $\xi(t)=[\xi_{1}^{T}(t),...,\xi_{N}^{T}(t)]^{T}$ and $\tilde{\sigma}(t)=\mathrm{diag}\{\sigma_{1}(t),$..., $\sigma_{N}(t)\}$. By Lemma IV.1, $\xi(t)$ can be partitioned into $\xi_{l}(t)=[\xi_{1}^{T}(t),\dots,$ $\xi_{M}^{T}(t)]^{T}$ and $\xi_{f}(t)=[\xi_{M+1}^{T}(t),\dots,$ $\xi_{N}^{T}(t)]^{T}$, $\vartheta(t)$ can be partitioned into $\vartheta_{l}(t)=$ $\vartheta_{1}^{T}(t),$ $\dots,\vartheta_{M}^{T}(t)]^{T}\text{and}$ $\vartheta_{f}(t)$ $\vartheta_{M+1}^{T}(t),\dots,\vartheta_{N}^{T}(t)]^{T}$, the adaptive gain $\tilde{C}(t)$ can be partitioned into $\tilde{C}_{l}(t)=\mathrm{diag}\{c_{1}(t),...,c_{M}(t)\}$ and $\tilde{C}_{f}(t)=\mathrm{diag}\{c_{M\!+\!1}(t),...,c_{N}(t)\!\},\text{the auxiliary time-varying gain}\ \tilde{\Sigma}(t)$ can be partitioned into $\tilde{\Sigma}_{l}(t)=\mathrm{diag}\{\Sigma_{1}(t),...,\Sigma_{M}(t)\}$ and $\tilde{\Sigma}_{f}(t)=\mathrm{diag}\{\Sigma_{M+1}(t),...,\Sigma_{N}(t)\}$, $\tilde{\sigma}(t)$ can be partitioned into $\tilde{\sigma}_{l}(t)=\mathrm{diag}\{\sigma_{1}(t),...,\sigma_{M}(t)\}$ and $\tilde{\sigma}_{f}(t)=\mathrm{diag}\{\sigma_{M+1}(t),...,\sigma_{N}(t)\}$, corresponding to the two subsets $\mathcal{V}_{l}$ and $\mathcal{V}_{f}$. Based on the closed-loop stochastic system (4), we obtain $d\vartheta_{l}(t)=(I_{M}\otimes A+(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t)\mathcal{L}_{11})\otimes B\mathcal{K})\vartheta_{l}(t)dt+(I_{M}\otimes C)\vartheta_{l}(t)dw(t)$, and $d\vartheta_{f}(t)=(I_{N-M}\otimes A+(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t)\mathcal{L}_{22})\otimes B\mathcal{K})\vartheta_{f}(t)dt+(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t)\mathcal{L}_{21}\otimes B\mathcal{K})\vartheta_{l}(t)dt+(I_{N-M}\otimes C)\vartheta_{f}(t)dw(t)$. Note that $\xi_{l}(t)=(\mathcal{L}_{11}\otimes I_{n})\vartheta_{l}(t)$ and $\xi_{f}(t)=(\mathcal{L}_{21}\otimes I_{n})\vartheta_{l}(t)+(\mathcal{L}_{22}\otimes I_{n})\vartheta_{f}(t)$. Then, we can get
@@
with $V_{l}(\xi(t),t)=\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds+$ $\eta_{2}^{l}\sum_{i=1}^{M}r_{i}(c_{i}(t)-\psi_{i})^{2}$ and $V_{f}(\xi(t),t)=\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}$ $c_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds+\eta_{2}^{f}\sum_{i=M+1}^{N}s_{i-M}(c_{i}(t)-\psi_{i})^{2}$, where $\sigma_{i}(t)=\xi_{i}^{T}(t)P\xi_{i}^{T}(t)$, $\eta_{1}^{l},\eta_{2}^{l},\eta_{1}^{f},\eta_{2}^{f}\geq 0$, $r_{i},s_{i}>0$, and $\psi_{i}$ are positive constants to be determined. According to the functional $\mathrm{It\hat{o}}$ formula and condition (I), it can be deduced that $dV(\xi(t),t)$ has the following form
$\begin{split}dV&(\xi(t),t)\!=\!\mathfrak{L}V(\xi(t),t)dt\!+\!\frac{\partial V(\xi(t),t)}{\partial\xi}\!(I_{N}\!\otimes\!C)\xi(t)dw(t),\end{split}$
where $\mathfrak{L}V(\xi(t),t)$ is defined as
$\begin{split}\mathfrak{L}V(\xi(t),t)=\mathfrak{L}V_{l}(\xi(t),t)+\mathfrak{L}V_{f}(\xi(t),t)\end{split}$ (12)
with
$\begin{split}\mathfrak{L}V_{l}(\xi(t),t)=&\eta_{1}^{l}\sum_{i=1}^{M}r_{i}\dot{c}_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds\\ &+2\eta_{2}^{l}\sum_{i=1}^{M}r_{i}(c_{i}(t)-\psi_{i})\dot{c}_{i}(t)\\ &+\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)\\ &-\int_{0}^{\sigma_{i}(t)}\mu k_{1}\frac{s}{c_{i}^{2}(t)}\dot{c}_{i}(t)(k_{2}+\frac{s}{{c}_{i}(t)})^{\mu-1}ds,\end{split}$
and
$\begin{split}\mathfrak{L}V_{f}(\xi(t),t)&=\eta_{1}^{f}\sum_{i=M+1}^{N}\!\!\!s_{i-M}\dot{c}_{i}(t)\!\!\int_{0}^{\sigma_{i}(t)}\!\!k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds\\ &+2\eta_{2}^{f}\sum_{i=M+1}^{N}s_{i-M}(c_{i}(t)-\psi_{i})\dot{c}_{i}(t)\\ &+\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)\\ &-\int_{0}^{\sigma_{i}(t)}\mu k_{1}\frac{s}{c_{i}^{2}(t)}\dot{c}_{i}(t)(k_{2}+\frac{s}{{c}_{i}(t)})^{\mu-1}ds.\end{split}$
Since $\mu,k_{1},k_{2}>0,\sigma_{i}(t)\geq 0,\ \text{we have}\ \mathfrak{L}V_{l}(\xi(t),t)\leq$ $\sum_{i=1}^{M}r_{i}\big[\eta_{1}^{l}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{l}c_{i}(t)-2\eta_{2}^{l}\psi_{i}\big]\dot{c}_{i}(t)+\eta_{1}^{l}\sum_{i=1}^{M}r_{i}$
@@
$c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)$, and $\mathfrak{L}V_{f}(\xi(t),t)\leq\sum_{i=M+1}^{N}s_{i-M}\big[\eta_{1}^{f}$ $\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{f}c_{i}(t)-2\eta_{2}^{f}\psi_{i}\big]\dot{c}_{i}(t)+\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)$ $\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)$. Let $\tilde{\psi}=\mathrm{diag}\{\psi_{1},...,\psi_{N}\}$. According to Lemma IV.1, $\tilde{\psi}$ can be partitioned into $\tilde{\psi}_{l}=\mathrm{diag}\{\psi_{1},...,\psi_{M}\}$ and $\tilde{\psi}_{f}=\mathrm{diag}\{\psi_{M+1},...,\psi_{N}\}$. Let $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$. Note that $\sum_{i=1}^{M}\sigma_{i}(t)=\xi_{l}^{T}(t)(I_{M}\otimes P)\xi_{l}(t)$ and $\xi_{l}(t)=(\mathcal{L}_{11}\otimes I_{n})\vartheta_{l}(t)$. By the definition of $\sigma_{i}(t)$, (10), (11) and using Lemma IV.1 and IV.2, we have
@@
$\begin{split}\mathfrak{L}&V(\xi(t),t)\\ \leq&\eta_{1}^{l}\xi_{l}^{T}\!(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &\!\!\!\!\!+\!\eta_{1}^{f}\xi_{f}^{T}\!(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &+\xi_{l}^{T}\!(t)\!\Big[\Phi_{t}^{l}\otimes\!\Gamma\Big]\!\xi_{l}(t)+\xi_{f}^{T}\!(t)\!\Big[\Phi_{t}^{f}\otimes\!\Gamma\Big]\!\xi_{f}(t),\end{split}$ (17)
@@
$\begin{split}\mathfrak{L}&V(\xi(t),t)\\ \leq&\eta_{1}^{l}\xi_{l}^{T}\!(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes(A^{T}P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &-\xi_{l}^{T}(t)\left[2\sqrt{\eta_{3}^{l}\check{\psi}_{l}}(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes\Gamma\right]\xi_{l}(t)\\ &+\eta_{1}^{f}\xi_{f}^{T}\!(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes(A^{T}P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &-\xi_{f}^{T}(t)\left[2\sqrt{\eta_{3}^{f}\check{\psi}_{f}}(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes\Gamma\right]\xi_{f}(t).\end{split}$
@@
$\begin{split}\mathfrak{L}V(\xi(t),t)\leq&-\eta_{1}^{l}\xi_{l}^{T}(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes I_{n}]\xi_{l}(t)\\ &-\eta_{1}^{f}\xi_{f}^{T}(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes I_{n}]\xi_{f}(t)\\ \leq&-\eta_{1}^{l}\varepsilon\xi_{l}^{T}(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes P]\xi_{l}(t)\\ &-\eta_{1}^{f}\varepsilon\xi_{f}^{T}(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes P]\xi_{f}(t)\\ =&-\eta_{1}^{l}\varepsilon\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\sigma_{i}(t)\\ &-\eta_{1}^{f}\varepsilon\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\sigma_{i}(t)\\ \leq&-\eta_{1}^{l}\varepsilon\sum_{i=1}^{M}r_{i}c_{i}(t)\!\!\int_{0}^{\sigma_{i}(t)}\!\!\!k_{1}\Big(k_{2}+\frac{s}{c_{i}(t)}\Big)^{\mu}ds\\ &-\eta_{1}^{f}\varepsilon\!\!\!\sum_{i=M+1}^{N}\!\!\!\!\!s_{i-M}c_{i}(t)\!\!\!\int_{0}^{\sigma_{i}(t)}\!\!\!\!\!\!k_{1}\Big(k_{2}\!+\!\frac{s}{c_{i}(t)}\Big)^{\mu}ds\\ \leq&-\varepsilon V(\xi(t),t)+\varphi(t),\end{split}$ (20)
@@
Step 2: By the functional $\mathrm{It\hat{o}}$ formula for $V(\xi(t),t)$ and the estimation (20) of $\mathfrak{L}V(\xi(t),t)\leq-\eta_{1}^{l}\varepsilon\xi_{l}^{T}(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes P]\xi_{l}(t)-\eta_{1}^{f}\varepsilon\xi_{f}^{T}(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes P]\xi_{f}(t)\leq-\eta_{1}\varepsilon\xi^{T}(t)[\tilde{S}(\tilde{C}(t)\tilde{\Sigma}(t))\otimes P]\xi(t)$ with $\eta_{1}=\eta_{1}^{l}\wedge\eta_{1}^{f}$ and $\tilde{S}=\mathrm{diag}\{R,S\}$, it can be deduced that
$\begin{split}&V(\xi(t),t)-V(\xi(0),0)\\ &=\int_{0}^{t}\mathfrak{L}V(\xi(s),s)ds+M_{v1}(t)\\ &\leq-\eta_{1}\varepsilon\!\int_{0}^{t}\!\xi^{T}(s)[\tilde{S}(\tilde{C}(s)\tilde{\Sigma}(s))\!\otimes\!P]\xi(s)ds+M_{v1}(t),\end{split}$ (21)
@@
$\begin{split}\lim_{t\to\infty}\sup V(\xi(t),t)<\infty\ a.s.\end{split}$
@@
$\begin{split}\ \int_{0}^{t}\!\mathrm{E}\eta_{1}\varepsilon\xi^{T}\!(s)[\tilde{S}(\tilde{C}(s)\tilde{\Sigma}(s))\!\otimes\!P]\xi(s)ds\leq\!\mathrm{E}V(\xi(0),0)\!<\!\infty,\end{split}$